Rachel has a 15km head start on Charlene. How long will it take Charlene to catch up with rachel if rachel travels at 70km per hour and charlene travels at 100 km per hour.
Answer by greenestamps(13195)



The rate at which Charlene catches up to Rachel is the difference between their two speeds, which is 100-70=30 km/h.

So the amount of time it take Charlene to make up 15km is 15/30 = 1/2 hour.